OpenStudy (anonymous):

given a point and the slope, write the equation in slope intercept form. (3,5) m=1/2

OpenStudy (anonymous):

First write it in point slope, and then convert it to slope-intercept: \[ (y-y_0)=m(x-x_0) \implies y-5=\frac{1}{2}(x-3) \]Now distribute and add 5 to both sides. Et voilá.

OpenStudy (anonymous):

is it correct to just plug the (x,Y) values into the slope intercept formula. and solve for b. in this case b=3.5 then just write it out y=1/2m+3.5 it comes out the same both ways but i dont know if its a fluke or what

OpenStudy (anonymous):

Yes, both ways are correct, my way is just slightly different. Do as you wish, or whichever you're more comfortable with.

OpenStudy (anonymous):

@precal

OpenStudy (precal):

haven't covered it yet, will next time....... but Lolwolf is correct

OpenStudy (precal):

btw the way you did it is the way I like to do it........lol